# Infra-red thermographic inversion in ST40

### v1

@@ Results
Following the guidelines of Sec. 3.3.2, time and space convergence are assessed. Table 1 and 2 respectively summarise the highlights of the two procedures, while Fig. 2 pictures the error behaviour.
@@
5.5E--04 60 4.0E--04 500 4.3E--05 3.07E--01
5.5E--04 60 1.0E--04 2000 1.0E--05 2.97E--01
5.5E--04 60 4.0E--05 5000 3.9E--06 2.95E--01
5.5E--04 60 2.0E--05 10000 1.7E--06 2.95E--01
5.5E--04 60 4.0E--06 50000 -- 2.94E--01
Table 2: Input parameters of the space convergence procedure (blue font for those which vary), and resulting errors. Reference FAHF simulation in azure (see text for details).
@@
2.3E--03 15 7.2E--06 27777 2.1E--03 1.10E--00
1.1E--03 30 7.2E--06 27777 4.0E--04 5.63E--01
5.5E--04 60 7.2E--06 27777 7.1E--05 2.52E--01
2.7E--04 120 7.2E--06 27777 2.5E--05 1.18E--01
1.4E--04 240 7.2E--06 27777 7.1E--06 5.72E--02
6.8E--05 480 7.2E--06 27777 -- 2.81E--02
In particular, Fig. 2a shows how the ideal first-order convergence (black dashed line) is closely achieved by the error points $\varepsilon_{*}(\Delta t)$. Similarly, Fig. 2b confirms the second-order convergence of the error in space $\varepsilon_{*}(\Delta x)$, hence verifying the space convergence of the algorithms. The minor departures of the data-points from the ideal dashed lines are not of concern, and could be ascribed to a combination of errors arising from interpolation and boundary conditions (Sec. 3.1.4).
In conclusion, the combination $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m (azure in Table 2) is chosen as a reference for further analyses (Sec. 4.2). This choice guarantees a satisfactory precision in $T_{i,j}^{n}$, at an affordable computational cost (158 s per sub-chord).
@@
According to Sec. 3.3.4, Fig. 3 depicts the behaviour of $\varepsilon_{*}(\mathcal{E})$ as a function of $\Delta x$, following from the data of which in Table 2 (last column).
The error $\varepsilon_{*}(\mathcal{E})$ is found to show a first-order convergence with $\Delta x=\Delta y$, and almost no dependence on $\Delta t$ (Table 1, last column). This is further investigated in Sec. 5.1. But because of the uncertainties customarily affecting experimental investigations of energy balance in tokamaks [51, 52], the error $\varepsilon_{*}(\mathcal{E})=5.72\%$ achieved according to Table 2 (azure) with the chosen $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m and $\Delta t=7.2\cdot 10^{-6}$ s (Sec. 4.1.1) is suitable for the analyses.
It is further noted that the actual sign [^2] Recovered by removing the absolute value at the numerator of Eq. (12). of $\varepsilon_{*}(\mathcal{E})$ is such that $\Delta\mathcal{E}_{\text{int}}(t)>\mathcal{E}_{\text{surf}}(t)\;\forall t\in[40;140]$ ms if $\Delta x\geq 5.5\cdot 10^{-4}$ m. Instead, no well-defined sign of the error is present in the high-resolution cases $\Delta x\leq 2.7\cdot 10^{-4}$ m, as the two energies converge together.
@@
The surface heat flux resulting from the set of simulations of Sec. 4.1.1 (hence with $\Delta x=\Delta y$) is analysed according to Sec. 3.1.6, at time instant $t_{*}=128$ ms.
As found for the energy error in Sec. 4.1.2, $q_{i}^{n}$ does also not show a dependence on $\Delta t$, with the Eich fit parameters varying by 2% at most. Therefore, only the results from the simulations of space convergence in Table 2 are detailed in Table 3. Full $q_{i}^{n}$ profiles in space are then shown by the dots in Fig. 4a, with their corresponding fitted Eich functions (solid lines). The main fitting parameters are pictured in Fig. 4b.
On the one hand, a strong dependence on $\Delta x=\Delta y$ is evidenced, with the main fitting parameters in fact changing by up to a factor 2.5 as the resolution increases. On the other hand, a convergence of both the heat flux profiles and the corresponding fitting parameters is successfully verified as $\Delta x=\Delta y$ drops. Any imperfect convergence is ascribed to the simplified fitting procedure here adopted (Sec. 3.1.6).
The present section thus justifies how $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m (Sec. 4.1.1) are also suitable choices from the perspective of $q_{i}^{n}$, which is in close reach of the converged value (azure in Table 3).
Table 3: Eich fitting parameters [6] and Eich integral (Eq. (6)) as a function of $\Delta x=\Delta y$ in the simulated cases of Table 2. Reference FAHF simulation in azure (see text for details).
@@
2.3E--03 15 2.3E--03 13 6.11E$+$06 5.35E--03 2.11E--03 2.53E$+$05 1.38E--03 3.28E$+$00 1.59E--01
1.1E--03 30 1.1E--03 26 9.16E$+$06 5.76E--03 1.70E--03 2.69E$+$05 2.72E--03 3.17E$+$00 2.23E--01
5.5E--04 60 5.5E--04 53 1.11E$+$07 6.68E--03 1.33E--03 2.45E$+$05 4.13E--03 2.87E$+$00 2.64E--01
2.7E--04 120 2.7E--04 107 1.19E$+$07 7.57E--03 9.48E--04 1.96E$+$05 4.98E--03 2.69E$+$00 2.83E--01
1.4E--04 240 1.4E--04 207 1.24E$+$07 1.04E--02 9.27E--04 3.50E$+$04 5.34E--03 2.10E$+$00 2.75E--01
6.8E--05 480 6.8E--05 426 1.26E$+$07 1.03E--02 7.81E--04 2.29E--08 5.52E--03 2.15E$+$00 2.78E--01
Figure 4: (a) Surface heat flux profiles computed by FAHF (dots), in the cases of Table 2 ($\Delta x=\Delta y$). Corresponding Eich fits in solid lines ($R^{2}>98.5\%$). (b) Main Eich fit parameters as a function of $\Delta x$.
Surface heat flux: $\Delta x\neq\Delta y$
The present paragraph revisits Sec. 4.1.3 in the light of $\Delta x\neq\Delta y$. For Eq. (5) depends on $\Delta y$ at the leading order, $\Delta x=5.5\cdot 10^{-4}$ m is here kept fixed, with $\Delta t=7.2\cdot 10^{-6}$ s (Table 2). Only $\Delta y$ is scanned according to Table 4 (first column), table which then reports the entire outcome of the scan at time $t_{*}=128$ ms. Heat flux profiles in space and main Eich fitting parameters are then depicted in Fig. 5.
The fitting parameters successfully converge towards fixed values, within a satisfactory 8% of their counterparts in Table 3 ($\Delta y=\Delta x$). This is however achieved at a diminished computational cost ($\propto N_{x}\cdot N_{y}\cdot N_{t}$): for instance, running FAHF with the $\Delta x=5.5\cdot 10^{-4}$ m and $\Delta y=6.3\cdot 10^{-5}$ m of Table 4 saves 45% of the computational time compared to the reference $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m, for the same $\Delta t=7.2\cdot 10^{-6}$ s. If $\Delta t$ was instead chosen according to Eq. (3), almost one order of magnitude worth of save in computational time could be achieved.
@@
5.5E--04 60 2.0E--03 14 6.37E$+$06 5.84E--03 1.99E--03 2.34E$+$05 1.76E--03 3.07E$+$00 1.63E--01
5.5E--04 60 1.0E--03 29 9.17E$+$06 6.08E--03 1.67E--03 2.53E$+$05 3.01E--03 3.03E$+$00 2.21E--01
5.5E--04 60 5.0E--04 58 1.08E$+$07 6.45E--03 1.20E--03 2.34E$+$05 4.28E--03 2.99E$+$00 2.57E--01
5.5E--04 60 2.5E--04 116 1.15E$+$07 8.21E--03 9.77E--04 1.84E$+$05 5.05E--03 2.49E$+$00 2.73E--01
5.5E--04 60 1.3E--04 232 1.19E$+$07 1.01E--02 8.73E--04 1.78E$+$04 5.37E--03 2.17E$+$00 2.63E--01
5.5E--04 60 6.3E--05 464 1.21E$+$07 1.02E--02 7.65E--04 2.57E--12 5.53E--03 2.16E$+$00 2.67E--01
5.5E--04 60 3.1E--05 928 1.22E$+$07 1.03E--02 7.21E--04 4.47E--09 5.60E--03 2.16E$+$00 2.71E--01
Figure 5: (a) Surface heat flux profiles computed by FAHF (dots), in the cases of Table 4 ($\Delta y\neq\Delta x$). Corresponding Eich fits in solid lines ($R^{2}>98.5\%$). (b) Main Eich fit parameters as a function of $\Delta y$.
@@
Downstream proving the choice of $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m as acceptable in terms of convergence errors (Sec. 4.1.1), energy balance error (Sec. 4.1.2), and heat flux behaviour (Sec. 4.1.3), the necessary condition for a well-posed cross-code comparison among FAHF and COMSOL is ensured.
Assessment of the geometrical error in FAHF
@@
According to Sec. 3.2, COMSOL is run in two (COMSOL2D) and three (COMSOL3D) dimensions, with temperature-independent material properties and with a single layer of molybdenum, hence matching FAHF models (Sec. 3.1.1). This allows insights in the geometrical error alone, as implied by Asm. #3.
@@
The close agreement between FAHF, COMSOL2D and COMSOL3D is depicted by Fig. 6, which shows the comparison of: (a) the error $\varepsilon^{n}(q)$ in time according to Eq. (10); (b) the error $\varepsilon_{i}(q)$ in space from Eq. (11); (c) the maximum heat flux in time, $\text{max}_{i\in[1;N_{x}]}\{q_{i}^{n}\}$; (d) the heat flux profiles in space at time instant $t_{*}=128$ ms. The errors are computed with the FAHF result always taken as the reference $\tilde{q}$, while (c) and (d) offer the possibility of qualitative insights.
@@
According to Fig. 6a, the satisfactory agreement between FAHF and COMSOL2D is evidenced for the entire time window $[40;140]\;\text{ms}$, with an error of $\varepsilon^{n}(q)=4.7\%\pm 1.3\%$. For FAHF vs. COMSOL3D, $\varepsilon^{n}(q)$ is clearly higher than its 2D counterpart, though still being an acceptable $8.3\%\pm 1.8\%$.
@@
A significant inaccuracy of FAHF is observed when the plasma: is started ($t\sim 10$ ms); reverts back to being limited ($t\sim 160$ ms, hence suddenly depositing its power elsewhere than the camera FoV on the divertor); is terminated ($t\sim 200$ ms), i.e. when evolving on a short timescale [^3] The maximum heat flux fluctuations in $[100;150]$ ms are a combination between: saturation of $T_{\text{IR}}$ (Sec. 2.3); considering the only maximum of heat flux; plasma vertical oscillation. These do not fall under the umbrella of ’fast dynamics’. (Fig. 6c). Accurately capturing fast transients would indeed require proper numerical handling [14], but it is not of concern in the present study.
@@
Fig. 6b suggests that similar conclusions hold for the error in space $\varepsilon_{i}(q)$, which satisfactorily attains values of $7.5\%\pm 1.4\%$ for FAHF vs. COMSOL2D all throughout the bulk of the tile, i.e. $x\in[2.5;30]$ mm. The enhanced error at the boundaries is imputable to FAHF there reverting to first-order BCs (Sec. 3.1.4). Slightly higher values of $10.0\%\pm 4.0\%$, though still acceptable, are found with respect to COMSOL3D.
Given the supportive evidence above, the appropriateness of adopting the multi-2D mode in FAHF in Asm. #3, rather than a full 3D treatment, is confirmed. This is also coherent with the findings of Ref. [46] concerning the multi-chord analysis.
@@
Figure 6: (a) Error $\varepsilon^{n}(q)$ in time. (b) Error $\varepsilon_{i}(q)$ in space. (c) Maximum heat flux in time $\text{max}_{i\in[1;N_{x}]}\{q_{i}^{n}\}$. (d) Heat flux profiles in space at time instant $t_{*}=128$ ms over the reference sub-chord.
@@
Starting from the COMSOL2D simulation of Sec. 4.2.2 with (i) temperature-independent material properties and a single layer of 29-mm molybdenum (Mo()($\cdot$)( ⋅ )), the set of simulations is enlarged to assess the model simplifications (Asms. #1 and #2) of Sec. 3.1.1. COMSOL is indeed run in 2D: (ii) with temperature-dependent material properties [53, 54, 55] and a single layer of 29-mm molybdenum (Mo$(T)$); (iii) with temperature-independent material properties and a double layer of 4-mm molybdenum and 25-mm copper-cromium-zirconium (Mo+CuCrZr()($\cdot$)( ⋅ )); (iv) with temperature-dependent material properties [56] and a double layer of 4-mm molybdenum and 25-mm copper-cromium-zirconium (Mo+CuCrZr$(T)$).
@@
In particular, $\varepsilon^{n}(q)<0.5\%\;\forall t\in[15;150]$ ms in Fig. 7a, and $\varepsilon_{i}(q)<1\%\;\forall x$ in Fig. 7b. This confirms that Asms. #1 and #2 are appropriate in the present case.
@@
Figure 7: (a) Error $\varepsilon^{n}(q)$ in time and (b) error $\varepsilon_{i}(q)$ in space, computed among the two-dimensional COMSOL cases with respect to the COMSOL simulation implementing FAHF-like models.

### v2

@@ Results
Any results discussed in this section is computed along CH#2 of Sec. 4.3.1.
@@
Following the guidelines of Sec. 4.3.2, time and space convergence are assessed. Tables 3 and 3 respectively summarise the highlights of the two procedures, while Fig. 3 pictures the error behaviour.
@@
5.5E--04 60 4.0E--04 500 4.3E--05 3.1E--01
5.5E--04 60 1.0E--04 2000 1.0E--05 3.0E--01
5.5E--04 60 4.0E--05 5000 3.9E--06 3.0E--01
5.5E--04 60 2.0E--05 10000 1.7E--06 3.0E--01
5.5E--04 60 4.0E--06 50000 -- 2.9E--01
Table 3: Input parameters of the space convergence procedure (blue font for those which vary), and resulting errors. Reference FAHF simulation in boldface.
@@
2.3E--03 15 7.2E--06 27777 2.1E--03 1.1E--00
1.1E--03 30 7.2E--06 27777 4.0E--04 5.6E--01
5.5E--04 60 7.2E--06 27777 7.1E--05 2.5E--01
2.7E--04 120 7.2E--06 27777 2.5E--05 1.2E--01
1.4E--04 240 7.2E--06 27777 7.1E--06 5.7E--02
6.8E--05 480 7.2E--06 27777 -- 2.8E--02
In particular, Fig. 3a shows how the ideal first-order convergence (black dashed line) is closely achieved by the error points $\varepsilon_{*}(\Delta t)$. Similarly, Fig. 3b confirms the second-order convergence of the error in space $\varepsilon_{*}(\Delta x)$, hence verifying the space convergence of the algorithms. The minor departures of the data-points from the ideal trend (e.g. $\Delta x=5.5\times 10^{-4}$ m in Fig. 3b) are not of concern, and could be ascribed to a combination of errors arising from interpolation and boundary conditions (Sec. 4.1.4).
In conclusion, the combination $\Delta t=7.2\times 10^{-6}$ s and $\Delta x=\Delta y=1.4\times 10^{-4}$ m (boldface in Table 3) is chosen as a reference for further analyses (Sec. 5.2). This choice guarantees a satisfactory precision in $T_{i,j}^{n}$, at an affordable computational cost of 158 s per sub-chord, scaling as $\propto N_{x}\times N_{y}\times N_{t}$. Notably, if the time step was instead chosen according to Eq. (7), i.e. $\Delta t=3\times 10^{-5}$ s, the computational time would drop to 35 s per sub-chord.
@@
Fig. 4 depicts the behaviour of $\varepsilon_{*}(\mathcal{E})$ as a function of $\Delta x$ (Sec. 4.3.4), following from the data of which in Table 3 (last column). The error $\varepsilon_{*}(\mathcal{E})$ is found to show a first-order convergence with $\Delta x=\Delta y$, and almost no dependence on $\Delta t$ (Table 3, last column). This is justified in Sec. 6.1.
The error $\varepsilon_{*}(\mathcal{E})=5.7\%$ achieved according to Table 3 (boldface) with the chosen $\Delta t=7.2\times 10^{-6}$ s and $\Delta x=\Delta y=1.4\times 10^{-4}$ m (Sec. 5.1.1) is suitable for the current analyses, also because of the uncertainties customarily affecting experimental investigations of energy balance in tokamaks [61, 62], at least of the same order of magnitude.
It is further noted that the actual sign [^4] Recovered by removing the absolute value at the numerator of Eq. (15). of $\varepsilon_{*}(\mathcal{E})$ is such that $\Delta\mathcal{E}_{\text{int}}(t)>\mathcal{E}_{\text{surf}}(t)\;\forall t\in[40;140]$ ms if $\Delta x\geq 5.5\times 10^{-4}$ m. Instead, no well-defined sign of the error is present in the high-resolution cases $\Delta x\leq 2.7\times 10^{-4}$ m.
@@
The surface heat flux resulting from the set of simulations of Sec. 5.1.1 (with $\Delta x=\Delta y$) is analysed according to Sec. 4.1.6, at time instant $t_{*}=128$ ms. As found for the energy error in Sec. 5.1.2, $q_{i}^{n}$ does also not show a dependence on $\Delta t$, with the Eich fit parameters varying by 2% at most. Therefore, only the results from the simulations of space convergence in Table 3 are detailed in Table 5. Full $q_{i}^{n}$ profiles in space are then shown by the dots in Fig. 5a, with their corresponding fitted Eich functions (solid lines). The main fitting parameters are pictured in Fig. 5c.
On the one hand, a strong dependence on $\Delta x=\Delta y$ is observed, with the main fitting parameters in fact changing by up to a factor 2.5 as the resolution increases. On the other hand, a convergence of both the heat flux profiles and the corresponding fitting parameters is successfully achieved as $\Delta x=\Delta y$ drops (magenta to cyan). Any imperfect convergence is ascribed to the simplified fitting procedure here adopted (Sec. 4.1.6).
The present section thus justifies how $\Delta t=7.2\times 10^{-6}$ s and $\Delta x=\Delta y=1.4\times 10^{-4}$ m (Sec. 5.1.1) are also suitable choices from the perspective of $q_{i}^{n}$, which is in close reach of the converged value (boldface in Table 5).
Surface heat flux: $\Delta x\neq\Delta y$
For Eq. (9) depends on $\Delta y$ to the leading order, $\Delta x=5.5\times 10^{-4}$ m is here kept fixed, with $\Delta t=7.2\times 10^{-6}$ s (Table 3). Only $\Delta y$ is scanned according to Table 5 (first column), table which then reports the entire outcome of the scan at time $t_{*}=128$ ms. Heat flux profiles in space and main Eich fitting parameters are then depicted in Fig. 5b and 5d, respectively.
The fitting parameters successfully converge towards fixed values, within a satisfactory 8% of their counterparts in Table 5 ($\Delta x=\Delta y$). This is however achieved at a diminished computational cost ($\propto N_{x}\times N_{y}\times N_{t}$): for instance, running FAHF with the $\Delta x=5.5\times 10^{-4}$ m and $\Delta y=1.3\times 10^{-4}$ m ($N_{x}=60$ and $N_{y}=232$) of Table 5 reduces the computational time of a factor $3.6$ compared to the reference $\Delta x=\Delta y=1.4\times 10^{-4}$ m ($N_{x}=240$ and $N_{y}=207$), for the same $\Delta t=7.2\times 10^{-6}$ s. If $\Delta t$ was instead independently chosen according to Eq. (7), $q_{0}$ and $\lambda_{q}$ matched within 5% could be achieved with a factor 6 worth of save in computational time: $\Delta t=9.1\times 10^{-5}$ s ($N_{t}=2223$) for $\Delta x=\Delta y=1.4\times 10^{-4}$ m, and $\Delta t=1.5\times 10^{-4}$ s ($N_{t}=1334$).
Table 4: Eich fitting parameters [40] and Eich integral (Eq. (3)) as a function of $\Delta x=\Delta y$ in the simulated cases of Table 3. Reference FAHF simulation in boldface (see text for details).
@@
2.3E--03 15 2.3E--03 13 6.11E$+$06 5.35E--03 2.11E--03 2.53E$+$05 1.38E--03 3.28E$+$00 1.59E$+$05
1.1E--03 30 1.1E--03 26 9.16E$+$06 5.76E--03 1.70E--03 2.69E$+$05 2.72E--03 3.17E$+$00 2.23E$+$05
5.5E--04 60 5.5E--04 53 1.11E$+$07 6.68E--03 1.33E--03 2.45E$+$05 4.13E--03 2.87E$+$00 2.64E$+$05
2.7E--04 120 2.7E--04 107 1.19E$+$07 7.57E--03 9.48E--04 1.96E$+$05 4.98E--03 2.69E$+$00 2.83E$+$05
1.4E--04 240 1.4E--04 207 1.24E$+$07 1.04E--02 9.27E--04 3.50E$+$04 5.34E--03 2.10E$+$00 2.75E$+$05
6.8E--05 480 6.8E--05 426 1.26E$+$07 1.03E--02 7.81E--04 2.29E--08 5.52E--03 2.15E$+$00 2.78E$+$05
@@
5.5E--04 60 2.0E--03 14 6.37E$+$06 5.84E--03 1.99E--03 2.34E$+$05 1.76E--03 3.07E$+$00 1.63E$+$05
5.5E--04 60 1.0E--03 29 9.17E$+$06 6.08E--03 1.67E--03 2.53E$+$05 3.01E--03 3.03E$+$00 2.21E$+$05
5.5E--04 60 5.0E--04 58 1.08E$+$07 6.45E--03 1.20E--03 2.34E$+$05 4.28E--03 2.99E$+$00 2.57E$+$05
5.5E--04 60 2.5E--04 116 1.15E$+$07 8.21E--03 9.77E--04 1.84E$+$05 5.05E--03 2.49E$+$00 2.73E$+$05
5.5E--04 60 1.3E--04 232 1.19E$+$07 1.01E--02 8.73E--04 1.78E$+$04 5.37E--03 2.17E$+$00 2.63E$+$05
5.5E--04 60 6.3E--05 464 1.21E$+$07 1.02E--02 7.65E--04 2.57E--12 5.53E--03 2.16E$+$00 2.67E$+$05
5.5E--04 60 3.1E--05 928 1.22E$+$07 1.03E--02 7.21E--04 4.47E--09 5.60E--03 2.16E$+$00 2.71E$+$05
Figure 5: Surface heat flux profiles computed by FAHF (dots), in the cases of (a) Table 3 with $\Delta x=\Delta y$ and (b) Table 5 with $\Delta y\neq\Delta x$. Corresponding Eich fits in solid lines ($R^{2}>98.5\%$). Main Eich fit parameters as a function of (c) $\Delta x$ and (d) $\Delta y$.
@@
Downstream proving the choice of $\Delta t=7.2\times 10^{-6}$ s and $\Delta x=\Delta y=1.4\times 10^{-4}$ m as acceptable in terms of convergence errors (Sec. 5.1.1), energy balance error (Sec. 5.1.2), and heat flux behaviour (Sec. 5.1.3), the necessary condition for a well-posed cross-code comparison of this case against COMSOL is ensured.
Assessment of the geometrical error in FAHF: CH#2 -- bulk of the tile
@@
According to Sec. 4.2.2, COMSOL is run in two dimensions (COMSOL2D) and three dimensions (COMSOL3D), with temperature-independent material properties and with a single layer of 29-mm molybdenum, hence matching FAHF’s Asms. #1 and #2. This allows insights in the geometrical error alone, as implied by Asm. #3. In the present section, simulation data are computed on CH#2 of Sec. 4.3.1, while Sec. 5.2.2 deals with the CH#1 instance.
@@
The satisfactory agreement between FAHF, COMSOL2D and COMSOL3D on CH#2 is pictured in Fig. 6, which shows the comparison of: (a) the maximum heat flux in time, $\text{max}_{i\in[1;N_{x}]}\{q_{i}^{n}\}$; (b) the heat flux profiles in space at time instant $t_{*}=128$ ms; (c) the error $\varepsilon^{n}(q)$ in time from Eq. (13); (d) the error $\varepsilon_{i}(q)$ in space from Eq. (14). In particular, (a) and (b) allow for useful qualitative insights, aided by the lack of camera saturation at $t_{*}=128$ ms along CH#2.
@@
According to Fig. 6c, the agreement between FAHF and COMSOL2D holds for the entire time window $[40;140]\;\text{ms}$, with an average error of $\varepsilon^{n}(q)=5\%$. For FAHF vs. COMSOL3D, $\varepsilon^{n}(q)$ is clearly higher than its 2D counterpart, though still being an acceptable $8\%$.
@@
A significant inaccuracy of FAHF is observed when the plasma: is started ($t\sim 10$ ms); reverts back to being limited on the centre column ($t\sim 160$ ms, hence suddenly depositing its power elsewhere than the divertor, Sec. 2.1); is terminated ($t\sim 200$ ms), i.e. when evolving on a short timescale [^5] The maximum heat flux fluctuations in $[100;150]$ ms are a combination between: saturation of $T_{\text{IR}}$ (Sec. 2.3); considering the only maximum heat flux; plasma vertical oscillation. These do not fall under the umbrella of ’fast dynamics’. (Fig. 6a). Accurately capturing fast transients would indeed require proper numerical handling [18], but it is not of concern in the present study, where the $\Delta t_{\text{IR}}$ (Table 1) would not be suited in the first place.
@@
Fig. 6d suggests that similar conclusions hold for the error in space $\varepsilon_{i}(q)$, which satisfactorily attains values of $8\%$ for FAHF vs. COMSOL2D all throughout the bulk of the tile, i.e. $x\in[2.5;30]$ mm. The enhanced error at the boundaries is imputable to FAHF there reverting to first-order BCs (Sec. 4.1.4). Slightly higher values of $10\%$, though still acceptable, are found with respect to COMSOL3D.
Given the supportive evidence above, the appropriateness of adopting a 2D approach in FAHF (Asm. #3), rather than a full 3D treatment, is confirmed for chords in the bulk of the tile, coherently with the findings of Ref.[48]. This conclusion especially holds during the time window of most interest, and its robustness is further enhanced by considering the implicit over-estimation of the error according to Sec. 4.3.3.
@@
Figure 6: CH#2 -- bulk of the tile: (a) maximum heat flux in time $\text{max}_{i\in[1;N_{x}]}\{q_{i}^{n}\}$; (b) heat flux profiles in space at time instant $t_{*}=128$ ms over the reference sub-chord; (c) error $\varepsilon^{n}(q)$ in time from Eq. (13); (d) error $\varepsilon_{i}(q)$ in space from Eq. (14). CH#1 -- vicinity of poloidal-running gap: (e) error $\varepsilon^{n}(q)$ in time from Eq. (13); (f) error $\varepsilon_{i}(q)$ in space from Eq. (14).
Assessment of the geometrical error in FAHF: CH#1 -- vicinity of poloidal-running gap
Conversely to Sec. 5.2.1, results on CH#1 are used in Fig. 6e and 6f. Only the errors are shown, for the sake of conciseness.
Average $\varepsilon^{n}(q)$ values of $7\%$ for FAHF vs. COMSOL2D and $22\%$ vs. COMSOL3D are found in $[40;140]$ ms on CH#1, hence above their counterparts on CH#2. The same holds for $\varepsilon_{i}(q)$ in space, with $11\%$ for the FAHF vs. COMSOL2D instance, and $36\%$ vs. COMSOL3D, above the values on CH#2.
This behaviour follows the expectations, for any poloidal-running gaps violates the assumption of axisymmetry along the toroidal direction leveraged by Asm. #3. The 2D approach is therefore shown to be more appropriate for chords away from the poloidal-running gaps ($\sim$ CH#2) than for chords in their proximity ($\sim$ CH#1).
This result should also caution on the selection of a representative region of interest, were a single-chord approach adopted (Sec. 4.1.2).
@@
Starting from the COMSOL2D simulation on CH#2 of Sec. 5.2.1 with (i) temperature-independent material properties and a single layer of 29-mm molybdenum (Mo()($\cdot$)( ⋅ )), the set of simulations is enlarged to assess the model simplifications in Asms. #1 and #2. COMSOL is indeed run in 2D on CH#2: (ii) with temperature-dependent material properties [63, 64, 65] and a single layer of 29-mm molybdenum (Mo$(T)$); (iii) with temperature-independent material properties and a double layer of 4-mm molybdenum and 25-mm copper-cromium-zirconium (Mo+CuCrZr()($\cdot$)( ⋅ )); (iv) with temperature-dependent material properties [66] and a double layer of 4-mm molybdenum and 25-mm copper-cromium-zirconium (Mo+CuCrZr$(T)$).
@@
In particular, $\varepsilon^{n}(q)<0.5\%\;\forall t\in[15;150]$ ms in Fig. 7a, and $\varepsilon_{i}(q)<1\%\;\forall x$ in Fig. 7b. This confirms that Asms. #1 and #2 are appropriate in the present case. Implementing more sophisticated models in FAHF could be accomplished by leveraging its modularity, were this be needed.
@@
Notably, the computational time of the COMSOL simulation Mo()($\cdot$)( ⋅ ), with settings of Sec. 4.2.2, is a factor 8 bigger than FAHF’s, notwithstanding the pre-processing time during the data handling, easily optimised in FAHF via Python tools [67].
Figure 7: (a) Error $\varepsilon^{n}(q)$ in time from from Eq. (13) and (b) error $\varepsilon_{i}(q)$ in space from Eq. (14), computed among the two-dimensional COMSOL cases with respect to the COMSOL simulation implementing FAHF-like models.